# A Practitioner’s Guide to Kolmogorov–Arnold Networks

### v1

@@ 7.4 Function decomposition
A complementary direction is given by Bühler et al. [204], who introduce KAN-SR, a symbolic regression framework that follows the recursive simplification strategies of Udrescu et al. [203]. Instead of fitting one large model, the method first checks whether the target function can be simplified by identifying separability (splitting variables into independent parts) or symmetry (invariance under transformations such as shifts or scalings). If such structure is found, the problem is broken down into smaller subproblems that are easier to solve. Each subproblem is then modeled with a compact KAN, and the results are combined to recover closed-form symbolic expressions. This divide–and–conquer strategy makes the learning process more interpretable and effective, especially in noisy or high-dimensional settings.
7.5 Hybrid/Ensemble & Data
A parallel stream of research improves accuracy not by new bases but by architectural integration—combining KANs with other neural modules. KANs have also been integrated with other neural components to form hybrid or ensemble models. These architectures combine the representation power of conventional neural networks with the interpretability and structured basis of KANs, aiming for higher accuracy, better conditioning, and improved use of data and physics priors.
Two representative examples highlight this direction:
  • He et al. [33] propose an MLP–KAN mixture-of-experts, where MLP blocks focus on representation learning while KANs specialize in function shaping. This division of labor improves flexibility without losing interpretability.
  • Xu et al. [50] introduce HPKM-PINN, a parallel hybrid in which KAN and MLP branches run side-by-side. Their outputs are fused through a learnable weight, yielding more accurate PDE solutions than either branch alone.
A more detailed hybrid design is KKAN, introduced by Toscano et al. [44]. This architecture integrates per-input feature learning with an explicit basis expansion. Each input coordinate $x_{p}$ is first mapped by a small MLP (the inner block) to one-dimensional features $\Psi_{p,q}(x_{p})$. A combination layer then sums across input dimensions to form $\xi_{q}=\sum_{p=1}^{d}\Psi_{p,q}(x_{p})$. For each $\xi_{q}$, an outer block applies a linear basis expansion
$G_{q}(\xi_{q})=\sum_{j=0}^{D}C_{q,j}\,b_{q,j}(\xi_{q}),$
where $b_{q,j}$ are simple basis functions (Chebyshev or Legendre polynomials, sine series [21], or RBFs). A final combination layer sums the outputs of all channels to produce the network output:
$\underbrace{f(x_{1},\dots,x_{d})}_{\text{output}}=\sum_{q=0}^{m}\underbrace{G_{q}\!\Big(\underbrace{\sum_{p=1}^{d}\Psi_{p,q}(x_{p})}_{\text{inner combination }(\xi_{q})}\Big)}_{\text{outer basis on }\xi_{q}}.$
In compact terms, KKAN follows the flow: inputs $\rightarrow$ inner MLPs $\rightarrow$ dimensionwise combination $\rightarrow$ basis expansion $\rightarrow$ final sum (see Fig. 19). The inner MLPs learn local, dimension-specific features, while the outer basis expansion provides a low-bias, well-conditioned representation. Toscano et al. [44] report that KKAN achieves lower test errors than both vanilla KANs and standard MLPs across tasks including function approximation, operator learning, and PDE solving.
Figure 19: KKAN architecture (after [44]). Each input coordinate $x_{p}$ These are summed across dimensions to form $\xi_{m}$, which are expanded via basis functions in the outer block. A final summation yields the output. Bases may include Chebyshev/Legendre polynomials, sine series, or RBFs. This “MLP inside + basis outside” layout mirrors the Kolmogorov–Arnold recipe.
7.6 Sequence/attention hybrids

### v2

@@ 7.8 Optimization & Adaptive Training
Bayesian and probabilistic methods. Probabilistic formulations are increasingly used to calibrate KANs or quantify prediction uncertainty. Lin [188] employed Bayesian optimization to tune RBF–KAN hyperparameters. Giroux and Fanelli [141] [^41] https://github.com/wmdataphys/Bayesian-HR-KAN introduced variational Bayesian inference for Higher–Order ReLU–KANs, placing priors on both weights and basis parameters; KL regularization improved robustness on stochastic Poisson and Helmholtz problems. Hassan [161] proposed a spline-based Bayesian–KAN, assigning Gaussian priors $c_{n}\!\sim\!\mathcal{N}(\mu_{n},\sigma_{n}^{2})$ to the activation coefficients for uncertainty quantification across layers.
A complementary direction is the probabilistic KAN of Polar and Poluektov [258] [^42] https://github.com/andrewpolar/pkan, which replaces Bayesian priors with an ensemble built by divisive data re-sorting. This approach learns multiple KANs over recursively refined subsets of the data, capturing full input-dependent output distributions—including multi-modality—while remaining computationally lightweight and highly parallelizable.
Initialization and depth-stabilized training. Beyond optimizer choice and probabilistic tuning, Rigas et al. [262] [^43] https://github.com/srigas/RGA-KANs show that deep Chebyshev-based KANs (cPIKANs) fail primarily due to variance blow-up across layers. They introduce a basis-agnostic Glorot-like initialization that preserves forward/backward variance and a Residual–Gated Adaptive block that stabilizes deep gradient flow. For spline KANs, Rigas et al. [276] [^44] https://github.com/srigas/KAN_Initialization_Schemes demonstrate that Glorot-style and power-law initializations substantially improve NTK conditioning and convergence across regression and PDE benchmarks, underscoring initialization as a key design lever for scalable KAN depth.
Advanced quasi-Newton methods. Kiyani et al. [263] show that self-scaled BFGS/Broyden optimizers markedly outperform Adam and standard BFGS on stiff PDEs. Their results reveal two effects: the PDE residual creates a highly “crinkled” loss landscape where first-order methods stall, and stronger optimizers cut this optimization error so effectively that smaller networks reach equal or better accuracy. One-liner: self-scaled quasi-Newton methods tame the PDE landscape and enable accurate, compact PIKANs. A related line of work by Poluektov and Polar  [264] [^45] https://github.com/andrewpolar/det5 replaces gradient training entirely with a Newton–Kaczmarz update on the spline/Gaussian coefficients, yielding a solver that is less sensitive to initialization and efficient for large datasets.
These works highlight the role of probabilistic reasoning for calibration, interpretability, and reliable prediction in settings with randomness or limited data.
8 Efficiency Improvement
Efficiency gains in KANs arise from two complementary directions: (i) exploiting hardware parallelism and compiler support, and (ii) reducing algebraic and basis-level complexity.
8.1 Parallelism, GPU, and JAX engineering
A major line of work accelerates KANs by making their core operations GPU-friendly. Iterative B-spline evaluations are replaced by non-iterative ReLU-power activations [48] or reformulated as CUDA-optimized matrix kernels [123, 133]. Model-level restructuring further boosts throughput: dual-matrix compression with Random Fourier Features [71], custom backward kernels to reduce memory stalls [193], and GPU-accelerated B-spline evaluation [124]. Architectural parallelism also plays a role, including multi-branch KAN–MLP hybrids [58] and domain-decomposition designs for PINNs [37, 122, 190]. At the framework level, JAX implementations [65, 118] leverage jit/XLA and optimized autodiff to deliver substantial end-to-end speedups.
8.2 Matrix optimization and parameter-efficient bases
A second strategy reduces computational cost by simplifying basis functions and matrix operations. Replacing spline bases with faster alternatives—ReLU-power [48, 123, 133], Chebyshev/Jacobi/Legendre polynomials [37, 197, 62, 63, 69], RBFs [188, 44], or wavelets [191]—lowers FLOPs and parameter counts while maintaining approximation quality. Additional savings come from structural compression: dual-matrix fusion with RFFs [71], sparsity-promoting regularizers [57], hierarchical knot refinement [46], and topology search via Differential Evolution [43]. Operator-aware designs further reduce cost by mixing spectral and spatial representations to keep matvecs near-diagonal [237], limiting mode mixing and memory traffic on variable-coefficient PDEs. Anant-Net [256] [^46] https://github.com/ParamIntelligence/Anant-Net extends these ideas with dimension-wise tensor-product sweeps and selective differentiation, both of which reduce high-dimensional autodiff complexity and are readily transferable to KANs. In a related direction, lookup multivariate KANs (lmKANs) replace univariate inner functions with multivariate spline lookup tables and CUDA-efficient evaluation, providing a parameter-rich yet computationally light alternative to linear layers [268] [^47] https://github.com/schwallergroup/lmkan.
Together, these techniques make KANs increasingly competitive in large- and high-dimensional settings by aligning basis, algebra, and implementation with modern hardware.
9 Sparsity & Regularization
Regularization plays a central role in improving the stability and generalization of KANs, especially in noisy, high-dimensional, or ill-posed regimes. Existing approaches fall into four broad families: sparsity penalties, Lipschitz/complexity control, weight decay, and implicit or dropout-style regularizers.
L1 sparsity and entropy balancing. Liu et al. [26] introduce an activation-level $\ell_{1}$ penalty,
$\|\varphi_{ij}\|_{1}=\tfrac{1}{N_{p}}\!\sum_{s=1}^{N_{p}}\!|\varphi_{ij}(x^{(s)})|,\qquad\|\Phi\|_{1}=\sum_{i,j}\|\varphi_{ij}\|_{1},$ (42)
supplemented by an entropy balancing term,
$S(\Phi)=-\sum_{i,j}\frac{\|\varphi_{ij}\|_{1}}{\|\Phi\|_{1}}\log\!\Big(\frac{\|\varphi_{ij}\|_{1}}{\|\Phi\|_{1}}\Big),$ (43)
leading to the regularized objective
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{pred}}+\lambda\big(\mu_{1}\!\sum_{\ell}\|\Phi_{\ell}\|_{1}+\mu_{2}\!\sum_{\ell}S(\Phi_{\ell})\big).$ (44)
Variants such as EfficientKAN [117] [^48] https://github.com/Blealtan/efficient-kan and DeepKAN [115] [^49] https://github.com/sidhu2690/Deep-KAN apply $\ell_{1}$ directly to weights; the same scheme appears in symbolic regression [53] and PDE learning [57]. For Neural ODEs, Koenig et al. [44] use a layerwise $\ell_{1}$ followed by pruning,
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{fit}}+\lambda_{\text{L1}}\|\theta\|_{1},$ (45)
eliminating redundant connections without accuracy loss.
Structured sparsity for symbolic regression. Bühler et al. [231] propose a composite regularizer for KAN-SR combining: (i) a magnitude penalty to suppress unused inputs, (ii) row/column entropies to localize attention, and (iii) an $\ell_{1}$ penalty on linear weights. This promotes compact, interpretable expressions while preventing the overspecification typical of symbolic regression networks.
Bayesian sparsity. Zou and Yan [243] introduce Probabilistic KANs (PKAN) and their sparse counterpart SSKAN, using Gaussian-process priors and spike–and–slab priors, respectively. These provide principled sparsity and uncertainty quantification, forming a Bayesian alternative to $\ell_{1}$-based pruning.
Lipschitz-driven regularization. Li et al. [274] show that KAN generalization is governed by a Lipschitz complexity depending on per-layer Lipschitz constants and mixed tensor norms. Their LipKAN architecture inserts 1-Lipschitz layers and replaces $\ell_{1}$ with a theoretically motivated $\ell_{1.5}$ penalty, providing tighter complexity control and improved generalization on complex distributions.
Low-rank RKHS regularization. Zhang and Zhou [273] establish generalization bounds for KANs whose activations lie in a low-rank RKHS (e.g., Wendland functions [275]), with complexity scaling in the activation rank rather than width. This motivates LoRA-style rank-constrained fine-tuning of pre-trained KAN activations: maintaining expressivity while regularizing model capacity.
L2 weight decay and extensions. Shen et al. [195] incorporate $\ell_{2}$ weight decay,
along with temporal smoothness and mutual-information terms,
to stabilize dynamics. Small $\ell_{2}$ penalties are similarly used in physics-informed KANs [37, 221] to prevent overfitting while preserving optimization stability.
Implicit and dropout-style regularizers. Implicit regularization arises when activations themselves bound outputs. Daryakenari et al. [65] show that nested activations such as $\tanh(\tanh(x))$ naturally confine outputs, smoothing gradients in Chebyshev-based KANs. Altarabichi et al. [224] propose DropKAN, which masks post-activation outputs rather than inputs, injecting noise directly into spline evaluations and offering more stable regularization than standard dropout.
10 Convergence and Scaling Laws
Classical approximation theorems guarantee that neural networks can approximate any continuous function, but provide little insight into convergence rates. Recent KAN studies fill this gap by deriving Sobolev-space error bounds, analyzing gradient-flow dynamics via NTK theory [225, 8], and documenting empirical scaling laws [187].
10.1 Theoretical Approximation Rates and Empirical Scaling Laws
Sobolev and Besov rates. Wang [53] proved that for $f\in W^{s}(\Omega)$, a KAN with depth $L$ and spline order $k$ satisfies
$\|f-g\|_{L^{p}(\Omega)}\;\leq\;C\,L^{-2s/d},$
corresponding to a parameter-rate $\mathcal{O}(P^{-2s/d})$, doubling the classical $\mathcal{O}(P^{-s/d})$ rate of standard neural and kernel approximators [226, 227]. Kratsios et al. [271] extend this to Besov spaces [272], showing that spline-based Res-KANs achieve the optimal Besov approximation rate on Lipschitz and fractal sets with dimension-free sample complexity. Complementing these approximation-centric results, Liu et al. [140] establish the first minimax statistical rates for KAN regression: spline-based additive and hybrid KANs achieve the optimal
$O\!\left(n^{-2r/(2r+1)}\right)$
for Sobolev-$r$ functions, together with an optimal knot-scaling rule $k_{n}\asymp n^{1/(2r+1)}$. These three perspectives jointly show that KANs match the best-known approximation and estimation rates of classical spline and sieve methods while retaining neural-network–style compositional structure.
Empirical scaling. Liu [26] observed that KAN RMSE follows a power law
$\ell\propto N^{-\alpha},$
with task-dependent $\alpha>0$. Increasing basis resolution (e.g., spline grid size $G$) consistently improves accuracy more than widening layers, echoing the theoretical depth benefit in [53] and the statistical spline-resolution rule in [140]. Together, these findings indicate that KANs achieve fast decay in both theory (Sobolev/Besov/minimax rates) and practice (empirical power laws), though [271] emphasizes that many of these gains reflect favorable optimization dynamics rather than asymptotic expressivity alone.
10.2 Spectral Bias and Frequency Learning Behavior
A model has spectral bias if low frequencies are learned before high ones [247, 248, 251]. In NTK terms, this corresponds to steep eigenvalue decay. Flatter spectra indicate more uniform frequency learning.
Reduced spectral bias in KANs. Wang [53] showed via Hessian–eigenvalue analysis that shallow KANs possess a well-conditioned landscape, enabling fast learning of oscillatory modes. Grid extension further widens the representable frequency band. Their experiments show KANs capturing high-frequency signals in a few iterations, unlike MLPs.
Farea [39] confirmed this in PDE settings: KANs and learnable-basis PINNs (e.g., Fourier/Jacobi) maintain flatter NTK eigenvalue spectra, indicating weaker spectral bias [252, 253]. However, they also showed that broader spectra can increase curvature, worsen conditioning, and destabilize training [250]. Wang [53] similarly noted that overly fine spline grids may amplify noise sensitivity, whereas modest coarsening can improve generalization.
Together, these results show that KANs learn high-frequency features more uniformly [53, 39], but at the cost of sharper loss landscapes—consistent with kernel-theoretic ill-conditioning phenomena [198, 199]. Basis choice must therefore balance spectral reach (e.g., Fourier-like) and stability (e.g., splines).
10.3 NTK-Based Convergence and Physics-Informed Extensions
Gradient-flow convergence. Gao [165] analyzed two-layer KANs under gradient flow, showing that the residual $s(t)$ evolves via
$\frac{d}{dt}s(t)=-G\,s(t),$
with $G=D^{\top}D$ the NTK Gram matrix. If $G\succ 0$, the loss decays exponentially:
giving a rigorous rate under appropriate learning rates. The same holds in expectation for SGD.
Physics-informed settings. For PDE residual losses,
$\mathcal{L}_{\mathrm{PDE}}(t)\;\leq\;\Big(1-\eta\,\tfrac{\lambda_{\min}}{2}\Big)^{t}\mathcal{L}_{\mathrm{PDE}}(0),$
but $\lambda_{\min}$ depends strongly on operator stiffness [165, 64]. Faroughi [64] showed that Chebyshev-based cKAN/cPIKAN models maintain better NTK conditioning and converge faster on diffusion, Helmholtz, Allen–Cahn, and beam PDEs. They also demonstrated that temporal decomposition and hybrid Adam–BFGS schedules improve conditioning, aligning with multi-stage results in [63, 62]. Initialization can also influence NTK conditioning: Rigas et al. [276] [^50] https://github.com/srigas/KAN_Initialization_Schemes show that variance-preserving Glorot-style and power-law schemes produce more stable initial NTKs for spline-KANs, leading to faster convergence.
10.4 Practical Trade-offs and Convergence Guarantees
KAN convergence is tightly controlled by the conditioning of the NTK Gram matrix: structured bases often enlarge the smallest eigenvalue $\lambda_{\min}$ relative to MLPs [165, 53], accelerating early optimization. Yet the same expressivity that broadens spectral coverage also increases curvature in the loss landscape, making training sensitive to basis resolution and learning-rate choices [39]. Effective practice therefore requires moderating basis richness to preserve numerical stability.
Physics-informed models highlight this balance. Chebyshev-KAN and cPIKAN architectures benefit from well-spread NTK spectra, producing faster PDE residual decay and more stable training [64, 63, 62]. But these gains diminish when grids are refined too aggressively or dimensionality grows, where operators become both costly to evaluate and harder to condition [42]. The result is a gradual trade-off between spectral reach and computational/numerical stability.
Generalization theory reaches the same conclusion from a different angle. Zhang and Zhou [273] show that KAN complexity is governed by $\ell_{1}$ coefficient norms and per-layer Lipschitz constants—precisely the terms that inflate when bases become overly sharp or high-frequency. Li et al. [274] formalize this perspective through Lipschitz complexity, deriving complexity-dependent generalization bounds and introducing LipKANs, which reduce Lipschitz growth via Lipschitz-controlled layers and $L^{1.5}$ regularization. Their results empirically demonstrate that controlling Lipschitz complexity mirrors the conditioning behavior seen in NTK dynamics: sharper bases yield higher curvature and poorer generalization.
Taken together, NTK convergence guarantees [165], approximation-rate theory [53, 273], Lipschitz-based generalization analyses [274], and empirical scaling laws [26] support a unified conclusion: KANs are expressive and parameter-efficient, but their practical performance depends on managing the expressivity–stability trade-off. The most effective models achieve wide spectral coverage while keeping curvature, conditioning, and coefficient growth under control [53, 39, 64].
11 Practical “Choose–Your–KAN” Guide
KANs offer a modular design space. The correct configuration depends on (i) the structure of the target function (smooth, oscillatory, discontinuous, periodic, multi-scale), (ii) computational constraints (time, memory, precision), and (iii) desired properties (accuracy, interpretability). Below is a clear, seven-step recipe synthesizing Sections 610.
Step 1 — Begin with a stable default. For general regression/classification, start with a cubic B-spline KAN on a uniform grid. This setup is mature, stable, and well-conditioned [26, 113]. For PDE or boundary-sensitive tasks, enable grid extension and, if needed, apply post-training refinement (Sec. 6.1; Fig. 1; [113]).
Step 2 — Select a basis, matched to the problem structure. Choose the basis family whose functional properties align with the anticipated solution:
  • Globally smooth / oscillatory: Chebyshev (ChebyKAN) with per-layer $\tanh$ normalization [196, 130, 49, 64].
  • Controllable smoothness or boundary shaping: Jacobi (fKAN, Jacobi-KAN) with fractional/domain parameters [127, 70].
  • Discontinuities, fronts, boundary layers: SincKAN (bandlimited) [49], rational rKAN [128], or DKAN (tanh-jump + smooth background) [60].
  • Periodic structure: FourierKAN or KAF (learnable RFF) [134, 71].
  • Local, multi-scale, or bursty features: Wavelets (Wav-KAN) [139, 191].
Step 3 — If speed matters, replace heavy components. For wall-clock efficiency, modify the basis or architecture:
  • Fast local bases: ReLU-KAN/HRKAN ($\mathrm{ReLU}^{m}$[123, 133]; or Gaussian RBF variants (FastKAN) [119].
  • Parallelism and compilers: JAX jit/XLA, fused CUDA kernels, multi-branch or domain-decomposition architectures [118, 65, 122].
  • Parameter economy: Prefer orthogonal polynomials or RBFs; prune using $\ell_{1}+$entropy penalties [37, 57, 26].
Step 4 — Stabilize via normalization, grid design, and Lipschitz control. Choose conditioning mechanisms suited to the basis, and reduce Lipschitz complexity where possible:
  • Chebyshev/Jacobi → apply $\tanh$ normalization to prevent blow-up of high-order modes [49, 65].
  • RBF/Wavelet → use min–max or sigmoid/tanh scaling to keep oscillatory activations bounded [231, 209].
  • Splines → employ extended grids or free-knot refinement near sharp gradients [46, 118].
  • All bases optionally insert $1$-Lipschitz “Lip” layers (tanh/sigmoid/ReLU) between $\Sigma$ and linear maps to reduce Lipschitz complexity and improve generalization [274].
Step 5 — Add physics and adaptive sampling only when needed. For PINNs/PIKANs, form
$\mathcal{L}=w_{u}\mathcal{L}_{\mathrm{data}}+w_{f}\mathcal{L}_{\mathrm{phys}}+w_{b}\mathcal{L}_{\mathrm{bc}},$
and enhance only when necessary:
  • Augmented Lagrangian or residual-based attention for conflicting constraints [1, 66, 52].
  • Residual-adaptive sampling (RAD) for error concentration [118].